OpenStudy (anonymous):

simplify each radical expression: radical 144x^3y^4z^5

OpenStudy (anonymous):

what is the square root of 144 ?

OpenStudy (anonymous):

we need to get that out of the way first this one is a bit harder than the last one, but only a bit harder, not much

OpenStudy (anonymous):

its 12

OpenStudy (anonymous):

ok good, so we know 12 comes outside of the radical now how many times does 2 go in to 3 ?

OpenStudy (anonymous):

just once

OpenStudy (anonymous):

yes, just once, but it leaves a remainder of 1, (unlike the last example) so \[\sqrt{x^3}=x^1\sqrt{x^1}=x\sqrt{x}\]

OpenStudy (anonymous):

we do the same for \(\sqrt{y^5}\) how many times does 2 go in to 5 ?

OpenStudy (anonymous):

2 times and a remainder of 1

OpenStudy (anonymous):

zactly so \[\sqrt{z^5}=z^2\sqrt{z}\]

OpenStudy (anonymous):

well it should be a \(y\) but you get the idea

OpenStudy (anonymous):

final answer is therefore \[12xy^2\sqrt{xy}\]

OpenStudy (anonymous):

ohh okay thank you so much i appreciate it

OpenStudy (anonymous):

yw